The financial statements for Castile Products, Inc., are given below: Castile Products, Inc. Balance Sheet December 31 Assets5. Book value per share. (Round your answer to 2 decimal places.) Book value per share
Answer #1

Ans:

Earning Per Share

= Net Income/No of Shares

=253960/(18000)

14.11

Dividend Payout Ratio

= Divined Per share/ Earning Per Share

=2.55/14.11

18.07%

Dividend Yield Ratio

= Dividend Per Share/ Market Value Per Share

=2.55/63

=4.05%

Price Earning Ratio

= Market Price Per Share/ Earning Per share

=63/14.11

4.46

Book Value Per Share

= Total Shareholder's Equity/ No of Shares

=774000/(180000/10)

43